# Quantum kinetic approach to the Schwinger production of scalar particles in an expanding universe

### v2

@@ 3 Physical observables
In this section we determine the main physical observables for the scalar field, namely the electric current and energy-momentum tensor, and express them in terms of kinetic functions $\mathcal{F}$, $\mathcal{G}$, and $\mathcal{H}$ introduced in the previous section.
@@ 3.1 Electric current
3.2 Energy-momentum tensor
The energy-momentum tensor of a scalar charged field $\chi$ can be found by varying the action with respect to the spacetime metric:
@@ 3.2 Energy-momentum tensor
Instead of pressure, $P_{\chi}=(1/3)g^{ij}T_{ij}^{\chi}$, it is more convenient to work with the trace of the energy-momentum tensor $T_{\chi}=g^{\mu\nu}T_{\mu\nu}^{\chi}$ which has the form:
@@ 4.1 Asymptotical behavior of the kinetic functions at high energies
$\mathcal{F}(\eta,\boldsymbol{p})$ $=\frac{e^{2}a^{4}(\boldsymbol{v}\cdot\boldsymbol{E})^{2}}{16\,\epsilon_{\boldsymbol{p}}^{4}}++\mathcal{O}(\epsilon_{\boldsymbol{p}}^{-5})\,,$ (37)
@@ 4.2 Electric current
$j^{i}=-\frac{e^{2}}{4}\int\frac{d^{3}\boldsymbol{p}}{(2\pi)^{3}}\,\frac{v^{i}v^{j}}{\epsilon_{\boldsymbol{p}}^{3}}\,\left(E^{j\prime}+2\frac{a^{\prime}}{a}E^{j}\right)+\mathcal{O}(\epsilon_{\boldsymbol{p}}^{-4})\,,$ (40)
where we wrote explicitly only the term which leads to the divergence. In order to extract the divergent contribution, we employ the dimensional regularization switching from the (3+1)-dimensional space time to a general $d$-dimensional one. Then, for the divergent integral in Eq. (40) by switching to the physical momentum $\boldsymbol{q}=\boldsymbol{p}/a$ we get:
@@
where $\Omega_{(d-1)}=2\pi^{(d-1)/2}/\Gamma(\tfrac{d-1}{2})$ is the full solid angle in $(d-1)$-dimensional space; $\mu_{\mathrm{r}}$ is a free parameter with the dimension of mass, which arises in order to compensate the mass dimension of the original integral and has a meaning of the renormalization energy scale (for simplicity, in what follows we set it $\mu_{\mathrm{r}}=m$); and $\gamma_{\mathrm{E}}\approx 0.577$ is the Euler-Mascheroni constant. Obviously, the first term in brackets in Eq. (41) tends to infinity in the limit $d\to 4$ which restores the correct dimension of our spacetime; therefore, it must be subtracted. However, together with this term, one can subtract any finite contributions. This ambiguity is resolved only by convention. One of the most popular subtraction schemes, the $\overline{\mathrm{MS}}$ scheme, requires to subtract the combination
@@
In the presence of this counterterm together with Eqs. (1) and (2) implies the following Maxwell equation for the electric field:
@@ 4.3 Energy density
The first term, $a_{1}$, renormalizes the cosmological constant, $a_{2}$ takes care of the Planck mass, and $a_{3}$ renormalizes the coefficient of the Starobinsky $R^{2}$ term [50]. The effective energy-momentum tensor which follows from such counterterms in the action can be obtained by varying Eq. (54) with respect to metric:
@@
which fully fix the counterterm in Eq. (54). Finally, subtracting from the integrand in Eq. (3.2) the terms shown in Eq. (4.3) and collecting the finite contributions in Eqs. (50)–(51), we obtain the expression for the regular part of the energy density:
@@
4.4 Trace of the stress-energy tensor
Finally, let us now consider the UV divergences in the trace of the energy-momentum tensor of the produced scalar particles. Expanding the integrand in Eq. (3.2) by using Eqs. (37)–(39), we collect the terms which might lead to the divergence:
@@ 4.4 Trace of the stress-energy tensor
It is straightforward to check that this expression is fully canceled by the trace of the effective energy-momentum tensor (4.3) if the coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by Eqs. (57)–(59). Thus, we need to add no new counterterms in order to cancel the divergence in the trace of the energy-momentum tensor. In order to get the finite part of $T_{\chi}$, we subtract from the integrand in Eq. (3.2) the terms in large-$p$ expansion up to $O(\epsilon_{\boldsymbol{p}}^{-3})$ inclusive, shown in Eq. (4.4) and take into account the finite contribution coming from the unity in Eq. (51). This leads to the following expression:
@@ 5 Conclusions
In this work, we describe the production of spinless charged particles (and antiparticles) in a homogeneous electric field in the expanding FLRW universe from the first principles. To this end, we study the particle excitations of a quantum scalar field which evolves on a classical time-dependent background consisting of a scale factor $a(\eta)$ and the electromagnetic vector-potential $\boldsymbol{A}(\eta)$. Because of the absence of static asymptotical states of this system in the infinite past and future we face a typical problem of how to define the vacuum state and particles in the expanding universe. Following the standard procedure [44, 45], we construct a physically reasonable approximation for the vacuum state – the adiabatic vacuum – in which the modes with very large momenta are almost not excited. However, this procedure is not unique as one can define adiabatic vacua of different adiabatic order and at different moments of time. In particular, in Ref. [43] and in the present work the same physical system was considered while difference is in the choice of adiabatic vacuum which leads to a slightly different results. We argue that the approach presented in this work leads to a system of equations which is more suitable for numerical simulations of the problem. Throughout the paper, we work with the conformal time in terms of which the FLRW metric is explicitly conformally flat. In a simple case of a conformally coupled massless scalar field in the absence of the electric field, the mode equation admits an exact solution which corresponds to the Bunch–Davies vacuum. Provided that initially the system was in this state, no particle creation occurs at any moment of time in future. This is a manifestation of a conformal symmetry of the physical system under consideration which makes its evolution in FLRW spacetime to be identical to that in the Minkowski spacetime. This symmetry was not respected in the system of equations in Ref. [43], where definition of the vacuum state is based on the Wentzel–Kramers–Brillouin solution to the mode equation written in physical time. In a general case, where the conformal symmetry is broken by the particle’s mass, nonminimal coupling $\xi\neq 1/6$, and/or the external electric field, the particle creation is described by the system of three quantum Vlasov equations for kinetic functions $\mathcal{F}(\eta,\boldsymbol{p})$, $\mathcal{G}(\eta,\boldsymbol{p})$, and $\mathcal{H}(\eta,\boldsymbol{p})$. The first one is the analog of a classical one-particle distribution function in the Boltzmann kinetic theory while the two others arise as auxiliary dynamical variables which describe the nonlocal in time process of particle creation. We study the asymptotical behavior of kinetic functions at large momenta and show that they decrease at large momenta much faster than in the case of Ref. [43]. Indeed, the leading term in the expansion in inverse powers of momentum is $\propto p^{-4}$ for $\mathcal{F}(\eta,\boldsymbol{p})$, $\propto p^{-3}$ for $\mathcal{G}(\eta,\boldsymbol{p})$, and $\propto p^{-2}$ for $\mathcal{H}(\eta,\boldsymbol{p})$ compared to $\propto p^{-2}$, $p^{-2}$, and $p^{-1}$ in the respective cases on Ref. [43]. This makes our new system of equations more suitable for numerical analysis on a lattice in the momentum space. The powerlike behavior of the kinetic functions at large momenta leads to the UV divergences in the main observables such as the electric current or the energy-momentum tensor of produced particles. We use the dimensional regularization in order to extract the divergent contributions and cancel them by introducing the counterterms to the action. The counterterms coincide with those derived in Ref. [43] and with a well-known results for the scalar quantum electrodynamics obtained by the conventional Green-function approach [49, 51, 44]. The obtained system of quantum Vlasov equations can be further used to describe the dynamics of particle production by a strong electric field in models of inflationary magnetogenesis. Another interesting problem which has a direct physical application within the Standard Model is to describe in a similar quantum kinetic approach the production of charged fermions. The presence of external magnetic field in addition to the electric one makes the dynamics much more interesting and may lead to chirality production via the chiral anomaly [52, 53]. The quantum kinetic description of particle production in non-orthogonal electric and magnetic fields is very relevant, e.g., for the axion inflation model; however, it is still missing in the literature. We plan to address these issues elsewhere. Acknowledgements O. O. S. is grateful to Prof. Kai Schmitz and all members of Particle Cosmology group for their kind hospitality at the University of Münster where the final part of this work was done.
@@
Funding The work was supported by the National Research Foundation of Ukraine (Project No. 2020.02/0062). The work of O. O. S. was sustained by a Philipp-Schwartz fellowship of the University of Münster. Conflicts of interest The authors declare no competing interests. Author Contributions All authors contributed in the writing and research of this paper. A. L. did computations and wrote the first draft of Sec. 2, 3 and Appendix A. O. S. performed computations in Sec. 4 and wrote Sec. 1, 4, and 5. Both authors have checked and approved the manuscript. Data availability No new data were created or analysed in this study.
@@ Appendix A Expansion in the inverse powers of momentum
Obviously, the operator $\hat{L}^{(0)}$ does not change the asymptotical behavior of a given term at large momenta while $\hat{L}^{(-1)}$ reduces the power of momentum by one. Further, we represent the kinetic functions as power series in inverse momentum: for $\mathcal{F}(\eta,\boldsymbol{p})$, $\mathcal{G}(\eta,\boldsymbol{p})$, and $\mathcal{H}(\eta,\boldsymbol{p})$ the series starts from the term $\propto p^{-4}$, $\propto p^{-3}$, and $\propto p^{-2}$, respectively. Then, substituting these expansions together with Eqs. (64)–(67) into the system of quantum Vlasov equations (2), we obtain the following set of equations:

### v3

@@ 3 Physical observables
In this section we determine the main physical observables for the scalar field, namely the electric current and energy–momentum tensor, and express them in terms of kinetic functions $\mathcal{F}$, $\mathcal{G}$, and $\mathcal{H}$ introduced in the previous section.
@@ 3.1 Electric current
3.2 Energy–momentum tensor
The energy–momentum tensor of a scalar charged field $\chi$ can be found by varying the action with respect to the spacetime metric:
@@ 3.2 Energy–momentum tensor
Instead of pressure, $P_{\chi}=(1/3)g^{ij}T_{ij}^{\chi}$, it is more convenient to work with the trace of the energy–momentum tensor $T_{\chi}=g^{\mu\nu}T_{\mu\nu}^{\chi}$ which has the form:
@@ 4.1 Asymptotical behavior of the kinetic functions at high energies
$\mathcal{F}(\eta,\boldsymbol{p})$ $=\frac{e^{2}a^{4}(\boldsymbol{v}\cdot\boldsymbol{E})^{2}}{16\,\epsilon_{\boldsymbol{p}}^{4}}+\mathcal{O}(\epsilon_{\boldsymbol{p}}^{-5})\,,$ (37)
@@ 4.2 Electric current
$j^{i}=-\frac{e^{2}}{4}\int\frac{d^{3}\boldsymbol{p}}{(2\pi)^{3}}\,\left[\frac{v^{i}v^{j}}{\epsilon_{\boldsymbol{p}}^{3}}\,\left(E^{j\prime}+2\frac{a^{\prime}}{a}E^{j}\right)+\mathcal{O}(\epsilon_{\boldsymbol{p}}^{-4})\right]\,,$ (40)
where we wrote explicitly only the term which leads to the divergence.
In order to extract the divergent contribution, we employ the dimensional regularization switching from the (3+1)-dimensional space time to a general $d$-dimensional one. Then, for the divergent integral in Eq. (40) by switching to the physical momentum $\boldsymbol{q}=\boldsymbol{p}/a$ we get:
@@
where $\Omega_{(d-1)}=2\pi^{(d-1)/2}/\Gamma(\tfrac{d-1}{2})$ is the full solid angle in $(d-1)$-dimensional space; $\mu_{\mathrm{r}}$ is a free parameter with the dimension of mass, which arises in order to compensate the mass dimension of the original integral and has a meaning of the renormalization energy scale (for simplicity, in what follows we set it $\mu_{\mathrm{r}}=m$); and $\gamma_{\mathrm{E}}\approx 0.577$ is the Euler–Mascheroni constant. Obviously, the first term in brackets in Eq. (41) tends to infinity in the limit $d\to 4$ which restores the correct dimension of our spacetime; therefore, it must be subtracted. However, together with this term, one can subtract any finite contributions. This ambiguity is resolved only by convention. One of the most popular subtraction schemes, the $\overline{\mathrm{MS}}$ scheme, requires to subtract the combination
@@
This counterterm together with Eqs. (1) and (2) implies the following Maxwell equation for the electric field:
@@ 4.3 Energy density
The first term, $a_{1}$, renormalizes the cosmological constant, $a_{2}$ takes care of the Planck mass, and $a_{3}$ renormalizes the coefficient of the Starobinsky $R^{2}$ term [52]. The effective energy–momentum tensor which follows from such counterterms in the action can be obtained by varying Eq. (54) with respect to metric:
@@
which fully fix the counterterm in Eq. (54).
Finally, subtracting from the integrand in Eq. (3.2) the terms shown in Eq. (4.3) and collecting the finite contributions in Eqs. (50)–(51), we obtain the expression for the regular part of the energy density:
@@
4.4 Trace of the energy–momentum tensor
Finally, let us now consider the UV divergences in the trace of the energy–momentum tensor of the produced scalar particles. Expanding the integrand in Eq. (3.2) by using Eqs. (37)–(39), we collect the terms which might lead to the divergence:
@@ 4.4 Trace of the energy–momentum tensor
It is straightforward to check that this expression is fully canceled by the trace of the effective energy–momentum tensor (4.3) if the coefficients $a_{1}$, $a_{2}$, and $a_{3}$ are given by Eqs. (57)–(59). Thus, we need to add no new counterterms in order to cancel the divergence in the trace of the energy–momentum tensor.
In order to get the finite part of $T_{\chi}$, we subtract from the integrand in Eq. (3.2) the terms in large-$p$ expansion up to $O(\epsilon_{\boldsymbol{p}}^{-3})$ inclusive, shown in Eq. (4.4) and take into account the finite contribution coming from the unity in Eq. (51). This leads to the following expression:
@@ 5 Conclusions
In this work, we describe the production of spinless charged particles (and antiparticles) in a homogeneous electric field in the expanding FLRW universe from the first principles. To this end, we study the particle excitations of a quantum scalar field which evolves on a classical time-dependent background consisting of a scale factor $a(\eta)$ and the electromagnetic vector-potential $\boldsymbol{A}(\eta)$. Because of the absence of static asymptotical states of this system in the infinite past and future we face a typical problem of how to define the vacuum state and particles in the expanding universe. Following the standard procedure [45, 46], we construct a physically reasonable approximation for the vacuum state—the adiabatic vacuum—in which the modes with very large momenta are almost not excited. However, this procedure is not unique as one can define adiabatic vacua of different adiabatic order and at different moments of time. In particular, in Ref. [44] and in the present work the same physical system was considered while difference is in the choice of adiabatic vacuum which leads to a slightly different results. We argue that the approach presented in this work leads to a system of equations which is more suitable for numerical simulations of the problem.
Throughout the paper, we work with the conformal time in terms of which the FLRW metric is explicitly conformally flat. In a simple case of a conformally coupled massless scalar field in the absence of the electric field, the mode equation admits an exact solution which corresponds to the Bunch–Davies vacuum. Provided that initially the system was in this state, no particle creation occurs at any moment of time in future. This is a manifestation of a conformal symmetry of the physical system under consideration which makes its evolution in FLRW spacetime to be identical to that in the Minkowski spacetime. This symmetry was not respected in the system of equations in Ref. [44], where definition of the vacuum state is based on the Wentzel–Kramers–Brillouin solution to the mode equation written in physical time.
In a general case, where the conformal symmetry is broken by the particle’s mass, nonminimal coupling $\xi\neq 1/6$, and/or the external electric field, the particle creation is described by the system of three quantum Vlasov equations for kinetic functions $\mathcal{F}(\eta,\boldsymbol{p})$, $\mathcal{G}(\eta,\boldsymbol{p})$, and $\mathcal{H}(\eta,\boldsymbol{p})$. The first one is the analog of a classical one-particle distribution function in the Boltzmann kinetic theory while the two others arise as auxiliary dynamical variables which describe the nonlocal in time process of particle creation.
We study the asymptotical behavior of kinetic functions at large momenta and show that they decrease at large momenta much faster than in the case of Ref. [44]. Indeed, the leading term in the expansion in inverse powers of momentum is $\propto p^{-4}$ for $\mathcal{F}(\eta,\boldsymbol{p})$, $\propto p^{-3}$ for $\mathcal{G}(\eta,\boldsymbol{p})$, and $\propto p^{-2}$ for $\mathcal{H}(\eta,\boldsymbol{p})$ compared to $\propto p^{-2}$, $p^{-2}$, and $p^{-1}$ in the respective cases on Ref. [44]. This makes our new system of equations more suitable for numerical analysis on a lattice in the momentum space.
The powerlike behavior of the kinetic functions at large momenta leads to the UV divergences in the main observables such as the electric current or the energy–momentum tensor of produced particles. We use the dimensional regularization in order to extract the divergent contributions and cancel them by introducing the counterterms to the action. The counterterms coincide with those derived in Ref. [44] and with a well-known results for the scalar quantum electrodynamics obtained by the conventional Green-function approach [51, 53, 45].
The obtained system of quantum Vlasov equations can be further used to describe the dynamics of particle production by a strong electric field in models of inflationary magnetogenesis. Another interesting problem which has a direct physical application within the Standard Model is to describe in a similar quantum kinetic approach the production of charged fermions. The presence of external magnetic field in addition to the electric one makes the dynamics much more interesting and may lead to chirality production via the chiral anomaly [54, 55]. The quantum kinetic description of particle production in nonorthogonal electric and magnetic fields is very relevant, e.g., for the axion inflation model; however, it is still missing in the literature. We plan to address these issues elsewhere.
Acknowledgements O. O. S. is grateful to Prof. Kai Schmitz and all members of Particle Cosmology group for their kind hospitality at the University of Münster where the final part of this work was done.
@@
Funding The work was supported by the National Research Foundation of Ukraine (Project No. 2020.02/0062). The work of O. O. S. was sustained by a Philipp-Schwartz fellowship of the University of Münster.
Conflicts of interest The authors declare no competing interests.
Author Contributions All authors contributed in the writing and research of this paper. A. L. did computations and wrote the first draft of Sec. 2, 3 and Appendix A. O. S. performed computations in Sec. 4 and wrote Sec. 1, 4, and 5. Both authors have checked and approved the manuscript.
Data availability No new data were created or analysed in this study.
@@ Appendix A Expansion in the inverse powers of momentum
Obviously, the operator $\hat{L}^{(0)}$ does not change the asymptotical behavior of a given term at large momenta while $\hat{L}^{(-1)}$ reduces the power of momentum by one.
Further, we represent the kinetic functions as power series in inverse momentum: for $\mathcal{F}(\eta,\boldsymbol{p})$, $\mathcal{G}(\eta,\boldsymbol{p})$, and $\mathcal{H}(\eta,\boldsymbol{p})$ the series starts from the term $\propto p^{-4}$, $\propto p^{-3}$, and $\propto p^{-2}$, respectively. Then, substituting these expansions together with Eqs. (64)–(67) into the system of quantum Vlasov equations (24), we obtain the following set of equations:
@@
Then, the terms $\mathcal{G}^{(-3)}$ and $\mathcal{G}^{(-4)}$ can be found from Eqs. (71) and (72), respectively: